OpenStudy (anonymous):

I NEEEEEEEEEED HELP PLEASE!!!!!!!!!!!!!!!!!!! PROBLEM BELOW!!!!!1

OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x – 6y = 12 9x + 2y = –24

OpenStudy (anonymous):

DOES ANYONE UNDERSTAND THIS?

OpenStudy (dape):

Have you done anything in on the problem?

OpenStudy (anonymous):

put the two equations in intercept form y=mx +b and graph; their point of intersection is the solution of the system

OpenStudy (anonymous):

NO

OpenStudy (dape):

Do you know what the graphing method of solving linear equations is?

OpenStudy (anonymous):